OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 1) and (1, 6)? A. -5 B.3/5 C.5/3 D. -7

OpenStudy (zzr0ck3r):

the slope of the line that passes through the points \((x_0,y_0)\) and \((x_1, y_1)\) is given by \(m=\frac{y_1-y_0}{x_1-x_0}\)

OpenStudy (anonymous):

Ok let me plug in the numbers like that brb

OpenStudy (anonymous):

@zzr0ck3r I keep getting 0.6, am I doing something wrong? :/

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@IMStuck can you help?

OpenStudy (zzr0ck3r):

\(\frac{6-1}{1-(-2)}=\frac{?}{?}\)

OpenStudy (anonymous):

6-1 = 5 and 1- (-2) =3

OpenStudy (imstuck):

formula is\[\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]Let's fill in your values.

OpenStudy (zzr0ck3r):

right 5/3

OpenStudy (anonymous):

Ok that makes sense

OpenStudy (imstuck):

\[\frac{ 6-1 }{ 1-(-2) }=\frac{ 5 }{ 3 }\]That's the slope. 5/3

OpenStudy (anonymous):

Thanks @zzr0ck3r and @IMStuck

OpenStudy (zzr0ck3r):

fo she z